# Automating High Quality RT Planning at Scale

### v2

@@ 2.1 Challenges of Clinical Plans for AI Training
OAR 3694 Brain DNU, Brain Push, Brachial Plexus, Brain sub, Brain2, Brainstem expd, Bstem_PRV, Cord + 0.5cm, Cord$<$ 45, Cord_PRV, Cord+3, Cord+5mm
Variations in contouring methods—such as differences in margins, anatomical boundaries, or slice intervals—introduce noise and bias, hindering the ability of models to generalize across different clinics or clinicians. Additionally, non-standardized planning preferences, including dose constraints and optimization strategies, lead to diverse and sometimes conflicting data inputs. As shown in Fig. 1, the number of clinically contoured OARs per patient varies significantly, complicating the normalization of plan quality to a consistent standard.
This lack of uniformity complicates the development of robust, generalized AI models that can learn effectively and be applied across multiple institutions. Additionally, much of the available RT data—whether from public sources (e.g., TCIA cohorts) or in-house clinical datasets—lacks clear, optimization objective definitions, making it difficult to align planned structures with the resulting RT dose maps. Table 1 highlights the heterogeneity in RT structure naming, which results in noisy matches between inputs and outputs, further complicating model training.
@@
Related Work. A closely related effort is the OpenKBP challenge [6]. Table 2 outlines the key differences between our approach and the OpenKBP data curation pipeline. Our work advances AI-based dose prediction by enhancing treatment planning across several dimensions. Specifically, we address more practical and diverse scenarios, including a broader range of treatment sites, variable prescribed doses tailored to individual patients, and an expanded set of ROI structures. Our dataset is also substantially larger, containing over ten times the number of plans compared to OpenKBP. Additionally, our plans are generated using one of the latest and most advanced planning systems—Eclipse 18.0.1 (released in 2023 by Varian Medical Systems). Unlike OpenKBP, our methodology aligns with current clinical guidelines for both planning objectives and quality assessment, ensuring greater relevance to real-world clinical practice.
[…]
Table 2: The comparison of our auto-planning and OpenKBP pipeline. The # patients and # plans are the numbers used in GDP-HMM challenge, the total number for our internal research topics is more.
Attribute OpenKBP [6] Ours Comments
Treatment Site HaN HaN & Lung
Prescribed Doses Fixed Variable OpenKBP change original clinical intent (PTV contours and prescription) to make data homogeneous. We keep the original clinical intent with the clinical prescription.
Number of PTVs Up to 3 Up to 3
Planning Tool CERR [2003] Eclipse [v.2023] Eclipse is from a leading Radiotherapy company Varian and widely used in clinical practice.
# Plans 340 3707
# Patients 340 1622
Quality Assessment No Yes We use the clinical Score card to evaluate the plan quality and remove or re-plan corner cases
# Structures up to 10/site up to 50/site Structures in our pipeline are Table 3 and 5.
2.2 Automatic High Quality Planning at Scale
Figure 2: Illustration of automated iterative RT planning (AIRTP) pipeline. The auto contouring, helper structures are executed with C++/python code. Curated data (CT and RTSTRUCT) are stored in DICOM format, and then imported into Eclipse system. Beam configuration, RapidPlan setup, photon optimization, dose calculation, quality assessment are conducted with Eclipse Scripting API.
The auto-planning pipeline consists of several steps, including auto-contouring, rule-based structure operations in order to create optimization helper structures, Eclipse engine as the main optimizer, scripting approach to iteratively run the optimization based on RapidPlan derived objectives, and planning quality assessment, as illustrated in Fig. 2.
Table 3: HaN RT structures for RapidPlan [28] and scorecard [2].
@@ 2.2 Automatic High Quality Planning at Scale
Table 6: HaN scorecard items
ROI Metric Type Input ROI Metric Type Input ROI Metric Type Input
PTVHighOPT VolAtDose 70.0 Gy PTVHighOPT DoseAtVol 99.5 % PTVHighOPT DoseAtVol 0.03 cc
PTVMid VolAtDose 63.0 Gy PTVMid DoseAtVol 99.5 % PTVMid-PTVHigh VolAtDose 66.2 Gy
PTV56 VolAtDose 56.0 Gy PTV56 DoseAtVol 99.5 % PTV56-PTV63 VolAtDose 58.8 Gy
SpinalCord05 DoseAtVol 0.03 cc SpinalCord05 VolAtDose 40.0 Gy SpinalCord05 VolAtDose 30.0 Gy
BrainStem03 DoseAtVol 0.03 cc Brain DoseAtVol 0.03 cc Brain VolAtDose 50.0 Gy
Pituitary MeanDose N/A Chiasm DoseAtVol 0.03 cc OpticNerveL DoseAtVol 0.03 cc
OpticNerveR DoseAtVol 0.03 cc LacrimalGlands MeanDose N/A CochleaR VolAtDose 40.0 Gy
CochleaL VolAtDose 40.0 Gy LensR DoseAtVol 0.03 cc LensL DoseAtVol 0.03 cc
EyeR DoseAtVol 0.03 cc EyeR MeanDose N/A EyeL DoseAtVol 0.03 cc
Mandible-PTV VolAtDose 70.0 Gy Mandible-PTV VolAtDose 60.0 Gy Mandible-PTV VolAtDose 50.0 Gy
Esophagus MeanDose N/A Esophagus DoseAtVol 0.03 cc OCavity-PTV MeanDose N/A
OCavity-PTV DoseAtVol 0.03 cc Larynx-PTV MeanDose N/A Thyroid-PTV MeanDose N/A
BrachPlexusL DoseAtVol 0.1 cc BrachPlexusR DoseAtVol 0.1 cc SubmandL-PTV MeanDose N/A
SubmandR-PTV MeanDose N/A TMJoint DoseAtVol 0.03 cc RingPTV70 DoseAtVol 0.03 cc
RingPTV63 DoseAtVol 0.03 cc RingPTV56 DoseAtVol 0.03 cc PosteriorNeck VolAtDose 35.0 Gy
Trachea MeanDose N/A Lungs VolAtDose 20.0 Gy Shoulders MeanDose N/A
Parotids-PTV MeanDose N/A Submand-PTV MeanDose N/A
Step 1: RT structures creation. As shown in Fig. 3, we first contour all the needed OARs based on AI Rad Companion of Siemens Healthineers [20]. According to [28], we create PTV helpers including ring structures and OARi-PTV, which are necessary as part of the RapidPlan [28] optimization dose objective derivation. The structures and the associated descriptions are shown in Table 3. All the structures are saved in DICOM format and are imported together with CT series, to the Eclipse workstation.
Step 2: Planning Script. There are three major modules in the planning script. Beam configuration includes setting up the fractions, beam geometries, and machine parameters. For each patient, we have both IMRT and VMAT setups following the same clinical goals. For the optimizer and dose calculation, we use TrueBeam with Photon Optimization version PO.18.0.1 and Dose Calculation AXB 18.0.1, respectively. To balance the time cost, the optimization convergence is set to “On” for VMAT.
For head-and-neck cancer treatment, we use four arcs for VMAT, alternating gantry rotations (CW and CCW) and setting collimator positions at 30° and 330°. For IMRT, we create two plans per patient using 9 and 15 evenly spaced angles. The objective definitions of the initial plan are primarily based on the RapidPlan model, following structure mapping. The RapidPlan model is obtained from the official Varian Medical Systems website [2].
For lung cancer treatment, we use three different angle ranges based on the tumor’s location, measured by the lateral distance from the isocenter to the patient’s mid-sagittal plane. If the distance exceeds 5 cm, we select either the left or right lung template, depending on the tumor’s side. If the distance is 5 cm or less, we choose angles that cover the full range. For IMRT, we use 7 fields (if lateral) or 9 angles, and for VMAT we use two arcs.
Step 3: Iterative Process. The initial plan is generated using RapidPlan [28] for head-and-neck treatment or mean/point objectives based on the scorecard [3] for lung treatment, as shown in Table 5. The iterative process is illustrated in Fig. 3(a). For each OAR, we extract DVH points from the current plan and apply a margin to set the objectives. Additionally, the metrics from the scorecard are included as objectives with some added margin. Evidence suggests that the plan quality will converge to a high score when the parameters are reasonably set. As shown in Fig. 3(b), we can reproduce a plan using only the 3D dose once the pipeline in Fig. 3(a) has converged.
More detailed steps are outlined in Algorithm 1, along with some module implementations shown in Figures 10 and 11 (Appendix).
Input: Patient data with CT, PTV/OAR contours, helper structures, prescribed doses for PTVs; ScoreCard for plan evaluation; beam configurations; RapidPlan for a head-and-neck scenario. Output: A series of plans until the ScoreCard convergence. 1:Initialize plan based on RapidPlan or ScorePlan items 2:for $n$ = 1 to MAX iteration do 3:   Initialize objective buffer $D$ 4:   for $s$ in PTV and its Ring structures do 5:      Set the upper / lower point objectives based prescribed dose 6:      Use the same priorities as in the initial plan 7:   end for 8:   Calculate the ScoreCard for current plan 9:   for $s$ in ScoreCard OARs and Helpers do 10:      Calculate the DVH Cumulative Data $f_{s}(\cdot)$ 11:      for $v$ in volume points {1, 10, 20, 30, 40, 60, 80do 12:         Calculate the dose $d=f_{s}(v)$ 13:         Calculate the margin and objective dose and priority based on 10. 14:         Add the objective to objective buffer. 15:      end for 16:      for items in ScoreCard do 17:         Calculate the margin and objective dose and priority based on 11. 18:         Add the objective to objective buffer. 19:      end for 20:   end for 21:   optimize with the current objective buffer. 22:   update current plan. 23:end for
Algorithm 1 Iterative Planning Processing
(a) iterative high-quality planning
(b) translate 3D dose prediction to deliverable plan
Figure 3: (a) shows the process that increase the quality of planning iteratively. (b) demonstrates the method translating predicted 3D dose to a deliverable plan. The way of defining objectives of above two pipelines are the same. The metrics in scorecard are shown in Table 6 for head-and-neck and Table 5 for lung cancer, respectively.
3 Planning and Results
3.1 Planning Environment
The first step in RT structure creation involves auto-contouring of OARs and the creation of helper structures, primarily based on clinical PTVs. The auto-contouring utilizes the latest version of AI-Rad Companion [20]. To accelerate this process, we developed a C++ executable that generates all the necessary structures with a harmonized naming convention and exports the results in DICOM format.
To run auto-planning at scale, we have built five GPU computing nodes in Miscorsoft Azure, in which instances of Eclipse 18.0.1 are installed. Each node has 64 GB Random-access memory (RAM) and 24 GB GPU memory. We use PyESAPI [4] and Python to import the DICOM input data and run the scripted algorithm to interact the Eclipse optimization engine. PyESAPI is a Python-based tool that enables rapid prototyping of C# based ESAPI built-in functionalities.
The public TCIA data sources (CT, PTVs, prescribed doses) that we utilized are from NSCLC-Cetuximab [10], NSCLC-Radiomics [1], Head-Neck Cetuximab [9], Head-Neck-PET-CT [31], HNC-IMRT-70-33 [11], HNSCC [18], HNSCC-3DCT-RT [8], TCGA-HNSC [37].
3.2 Plan Quality Analysis
Figure 4: Examples of head-and-neck and lung plans. In the GDP-HMM challenge, each lung cancer patient has two plans: one IMRT and one VMAT; each head-and-neck cancer patient has three plans: two IMRT and one VMAT.
Fig. 4 demonstrates plan examples for lung cancer and head-and-neck cancer patients. Our AIRTP can generate arbitrarily large number of plans with varying parameters, given sufficient time and computing resources. As the first public release of our data, we include two plans per patient for the lung site and three plans per patient for the head-and-neck site. Additional examples can be found in Appendices F and G.
Fig. 5 demonstrates that our iterative refinement improves the plan quality of a strong RapidPlan baseline [28]. The plan quality metrics follow the a modified scorecards from [3] for lung and [2] for head-and-neck treatments, where large negative values are included when criteria are not met.
(a) Score improvement of IMRT plans.
(b) Score improvement of VMAT plans.
Figure 5: Quality improvement achieved through iterative refinement across head-and-neck cohorts.
3.3 Comparison with Clinical Plans
Two randomly selected head-and-neck cases from different cohorts are illustrated in Fig. 6. One notable observation is that the AIRTP plan demonstrates improved PTV homogeneity without compromising OAR sparing. However, it is important to emphasize that we cannot definitively claim that the AIRTP plan is superior to the clinical manual plan based solely on DVHs or scorecards, as the planning process is inherently subjective and clinical plans may follow different protocols. On a positive note, we observed no significant quality drop compared to clinical plans. This fully automated AIRTP process eliminates human bias, paving the way for unbiased large-scale AI training.
3.4 Plan Reproducibility Given 3D Dose
The advantages of high plan reproducibility are twofold. First, when a dose prediction model is properly trained, it enables the rapid generation of visualized 3D dose distributions and dose-volume histograms (DVHs) based on patient data and beam configuration. This allows users to quickly assess and adjust the treatment plan. For example, a typical head-and-neck case requires approximately 20 minutes for a single forward-planning session, whereas deep learning-based dose prediction methods, such as those described in [14], can generate results in just 0.2 seconds (excluding data loading and preprocessing time). Second, this high reproducibility provides an opportunity to evaluate dose predictions in the context of downstream clinical tasks.
(a) HaN: reference (solid) vs. reproduce (dash).
(b) Lung: reference (solid) vs. reproduce (dash).
3.5 Beyond the Topics of Generalizable Dose Prediction
Except for dose prediction, researchers have explored AI applications including fluence prediction [33], leaf sequencing [13], dose computation [35]. With our pipeline, we can create large-scale training data for all these modules. As in Fig. 8, we show a few examples of fluence maps, leaf sequences, and dose volumes generated from our pipeline.
Figure 8: Data examples (screenshots from Eclipse visualization) are provided to support AI training for tasks such as leaf sequencing, fluence prediction, dose calculation, and/or their combinations.
3.6 AI benchmarking for Generalizable Dose Prediction
Table 7: Data splits (train/valid/test).
# Plans # Patients
HaN 1428 / 224 / 298 499 / 77 / 100
Lung 1450 / 132 / 200 776 / 70 / 100
Total 3732 1622
Table 8: Results of different phases.
MAE ($\downarrow$) Plan Quality Score ($\uparrow$)
Sanity Check 1.981 103.09
Validation 2.540 110.76
Testing 3.454 118.78
Many backbones used in dose prediction have adopted architectures originally designed for medical image segmentation [14]. In the third-party medical image segmentation benchmark TouchStone [7], which includes a test set of 6,933 CT volumes, and the nnUNet revisit benchmark [22], which evaluates six public datasets, MedNeXt [30] achieves top performance, significantly outperforming popular Transformer-based backbones by a large margin (up to a 25% DSC improvement). Accordingly, we adopt MedNeXt as an educational baseline for generalizable dose prediction using publicly available data from our AIRTP, as illustrated in Fig. 13.
Inspired by [14], we introduce the angle plate and beam plate to encode spatial information related to beam angles. Additionally, categorical information (e.g., treatment mode, anatomical site) is incorporated as an extra input channel. The training dataset is divided into dev-train and dev-valid, where model weights are updated using dev-train, and the best checkpoint is selected based on dev-valid.
To facilitate participation, we provide a complete implementation—including code, data, tutorials, and a pre-trained model—on GitHub, enabling users with minimal radiotherapy (RT) background to easily engage with the challenge. An overview of the data distribution is presented in Table 8, while the baseline results are summarized in Table 8.333The plan quality scores for validation / testing phases are based on a randomly selected subset of 50 plans.
4 Discussion
Conclusion. In this work, we propose an automated RT planning pipeline that includes OAR auto-contouring, scripted RT configuration and objective setup, scripted optimization and calculation, as well as automated iterative refinement guided by quality evaluation. The results from our pipeline have been used in the GDP-HMM challenge, one of few challenges selected by the AAPM community. Our proposed method, along with the released dataset, has the potential to advance the field of artificial intelligence in radiotherapy.
As an initial step toward automating high-quality planning at scale, our work has several limitations compared to expert planning in clinical practice. First, clinical planning involves detailed patient-specific considerations, such as variations in the contouring of PTVs/OARs, auxiliary structures, and various dose constraints. In our pipeline, we focus primarily on clinical PTVs and prescribed doses as the key patient-specific factors, simplifying the process due to the lack of access to more detailed patient data. Second, different institutions often follow different protocols, and even within the same institution, planners may have individual preferences. Our pipeline does not currently account for such subjective variations. While these preferences are partially reflected in the scorecards, we use only one standardized scorecard per treatment site. A future direction is to collaborate with dosimetrists to develop multiple scorecards that better capture diverse clinical preferences and practices. Third, unlike more well-established research fields, our comparative analyses of iterative refinement and re-planning strategies are limited by the lack of directly relevant literature. Although we made efforts to compare the DVHs of AIRTP plans with clinically approved plans, this comparison may not be entirely fair, as the AIRTP and clinical plans follow different guidelines. Fourth, this paper focuses on 1) novel methodology to automate high-quality planning at scale, 2) analyses of planning results, 3) open sourcing high-quality data/tutorial/code/pre-trained model for a AAPM challenge and other RT research topics. the summary and clinical impact analysis of the GDP-HMM challenge are beyond the scope of this paper and will be provided following the conclusion of the challenge.
Outlook. Compared to other AI domains such as natural language processing and computer vision, the availability of large-scale training data in medical imaging is often limited. In radiotherapy, this challenge is further amplified by the complexity, subjectivity, noise, and limited scale of clinical data, as discussed in Section 2.1. Despite these obstacles, our work represents a significant step toward enabling large-scale AI research in radiotherapy. Notably, our pipeline is designed to be flexible, making it possible to simulate data at scale with sufficient computational resources and time. We envision this pipeline as a foundational tool for advancing AI applications in radiotherapy and facilitating future research in this critical field with more required data.
Appendix
[…]
Figure 9: An illustration of what a huge gap between AI models in RT planning and some well-known successful AI models in other domain. The y-axis is in base-10 logarithmic scale.
We add more descriptions about the models in Fig. 9:
  • Llama 3 [15]: An advanced large language model (LLM) designed for natural language understanding and generation, offering improved efficiency and performance over its predecessors.
  • Stable Diffusion [29]: A state-of-the-art latent diffusion model for generating high-quality images from text prompts, widely used for AI-driven image synthesis.
  • Segment Anything Models (SAM) [24]: A foundational vision model capable of segmenting objects in images with minimal user input, enabling versatile and efficient image annotation.
  • Medical SAM (Med-SAM) [27]: A specialized adaptation of SAM for medical imaging, fine-tuned to segment anatomical structures with high accuracy across diverse modalities.
  • RETFound [36]: A foundation model for retinal imaging that leverages self-supervised learning to enhance disease detection and diagnosis in ophthalmology.
  • Medical AI for Synthetic Imaging (MAISI) [19]: A generative AI framework designed to synthesize medical images for training and validation, improving AI model generalization and robustness.
  • Dose Prediction Models [14, 6, 12, 23]: AI-based models for predicting radiation therapy dose distributions, as the task in GDP-HMM challenge.
  • Fluence Prediction Models [33, 26]: Computational models that predict fluence maps in radiation therapy, helping to optimize beam intensity for effective dose delivery.
  • Leaf Sequencing Models [13, 21]: Algorithms that translate optimized fluence maps into deliverable multi-leaf collimator (MLC) sequences, ensuring efficient and accurate radiation therapy delivery.
Many of the well-known successful AI models have been trained with large-scale and well-curated data. In contrast, there are significant challenges for AI application in RT: 1) AI training samples in the RT planning domain are rather limited, usally under 1,000 sample; 2) missing and noisy information usually exists during clinical RT data collection; and 3) most well-curated RT training samples are not shared publicly. With the enthusiasm to advance RT planning with AI, we conducted this study to address those challenges.
Appendix B Code Blocks for Algorithm 1
@@
Appendix C More about GDP-HMM Challenge
C.1 Data Distribution
@@
C.2 Baseline Method
As shown in Fig. 13, the educational baseline applies the MedNeXt as the backbone; due to its superior performance in multiple large-scale segmentation benchmarks. Fundamentally, the dose prediction is an conditional image generation problem. The CT images, PTV & OAR masks, planning mode, treatment site, and beam geometries etc., are conditions for 3D dose prediction. We follow the strategy of from [14], concatenating all the conditions as multi-channel 3D input. The loss function used in the baseline is pretty simple: the mean absolute error (MAE) of the predicted dose and the reference dose.
@@ C.2 Baseline Method
C.3 Result Samples from the Baseline
@@
We present four samples from our validation cohort, illustrating distinct scenarios that a universal AI model should be capable of handling. Figures 14 and 15 depict treatment plans for the same head-and-neck cancer patient using different modalities: IMRT and VMAT. Similarly, Figures 16 and 17 show treatment plans for the same lung cancer patient, also with IMRT and VMAT. Each case includes a comparison of the predicted and reference 3D dose distributions, along with corresponding dose-volume histograms (DVHs).
@@
Appendix D Scope of This Paper and the Planned Challenge Summary Paper
A summary paper for the challenge is planned to be published several months after its conclusion. It is important to highlight the distinctions between this paper and the upcoming challenge summary paper.
This paper presents our key contributions, including: 1) A novel AIRTP pipeline for automatically generating high-quality radiotherapy plans. 2) A comprehensive dataset (DICOM and NPZ format) and accompanying toolkit to facilitate research. 3) Baselines and tutorials for the challenge.
The advantages of AIRTP, outlined in Section 1, enable large-scale AI training in radiotherapy, addressing the limitations of previous studies, which often relied on fewer than 1,000 plans. Our pipeline streamlines data generation, allowing for significantly larger training scales.
Additionally, our dataset supports a broad range of research topics beyond dose prediction, the primary focus of the challenge. We also provide a toolkit for reading and visualizing raw DICOM data, enhancing usability for researchers.
Furthermore, we establish a benchmark for the GDP-HMM challenge using the state-of-the-art MedNeXt model. We provide detailed tutorials, open-source code, pre-trained model weights, and thorough result analyses, offering participants and future researchers a strong foundation for further exploration.
The challenge itself is focused solely on dose prediction, which is just one of many critical topics in radiotherapy. The planned challenge summary paper will primarily cover: 1) A comparative analysis of participants’ methods, evaluating AI techniques and performance results. 2) A deeper discussion on how AI can enhance radiotherapy planning through 3D dose prediction, with a focus on clinical implications.
This planned summary will provide insights into the practical impact of AI in radiotherapy, complementing the technical contributions outlined in this paper.
Appendix E Detailed Scorecard Examples
Table 10: An example of 9-angle IMRT for head-and-neck cancer.
StructId MetricType Input Output (Initial) Score (Initial) Output (AIRTP) Score (AIRTP)
PTVHighOPT VolumeAtDose 70.0 Gy 97.00 % 19.00 97.00 % 19.00
PTVHighOPT DoseAtVolume 99.5 % 69.07 Gy 1.33 69.10 Gy 1.34
PTVHighOPT DoseAtVolume 0.03 CC 74.83 Gy 8.55 74.50 Gy 8.78
PTVLow VolumeAtDose 56.0 Gy 99.39 % 14.90 99.36 % 14.89
PTVLow DoseAtVolume 99.5 % 55.82 Gy 1.46 55.85 Gy 1.47
PTVLow-PTVHigh VolumeAtDose 58.8 Gy 75.89 % 0.82 60.61 % 3.88
SpinalCord_05 DoseAtVolume 0.03 CC 45.44 Gy 5.82 46.85 Gy 5.26
SpinalCord_05 VolumeAtDose 40.0 Gy 5.42 % 1.37 3.92 % 1.61
SpinalCord_05 VolumeAtDose 30.0 Gy 45.87 % -0.44 33.38 % 1.16
BrainStem_03 DoseAtVolume 0.03 CC 49.56 Gy 3.01 45.50 Gy 3.11
Brain DoseAtVolume 0.03 CC 57.77 Gy 1.61 54.10 Gy 1.64
Brain VolumeAtDose 50.0 Gy 0.64 CC 3.00 0.08 CC 3.00
Pituitary MeanDose None 5.50 Gy 0.95 5.74 Gy 0.95
OpticNerve_L DoseAtVolume 0.03 CC 3.55 Gy 2.94 3.68 Gy 2.93
OpticNerve_R DoseAtVolume 0.03 CC 3.37 Gy 2.94 3.72 Gy 2.93
Cochlea_R VolumeAtDose 40.0 Gy 0.00 % 3.00 0.00 % 3.00
Cochlea_L VolumeAtDose 40.0 Gy 0.00 % 3.00 0.00 % 3.00
Lens_R DoseAtVolume 0.03 CC 1.75 Gy 2.33 1.83 Gy 2.32
Lips MeanDose None 24.83 Gy 2.58 20.63 Gy 4.68
Mandible-PTV VolumeAtDose 70.0 Gy 0.00 % 5.00 0.00 % 5.00
Mandible-PTV VolumeAtDose 60.0 Gy 0.16 % 2.00 0.01 % 2.00
Mandible-PTV VolumeAtDose 50.0 Gy 16.75 % 1.78 10.52 % 1.86
Esophagus MeanDose None 17.51 Gy 3.62 19.06 Gy 3.55
Esophagus DoseAtVolume 0.03 CC 54.07 Gy 3.05 56.14 Gy 2.98
OCavity-PTV MeanDose None 39.58 Gy 1.67 33.19 Gy 3.63
OCavity-PTV DoseAtVolume 0.03 CC 59.82 Gy 1.76 57.35 Gy 1.92
RingPTVHigh DoseAtVolume 0.03 CC 72.24 Gy 2.70 71.70 Gy 3.86
RingPTVLow DoseAtVolume 0.03 CC 63.92 Gy -6.82 63.11 Gy -5.74
Posterior_Neck VolumeAtDose 35.0 Gy 52.65 % -2.12 45.78 % 0.68
Trachea MeanDose None 13.16 Gy 2.11 13.66 Gy 2.09
Lungs VolumeAtDose 20.0 Gy 13.36 CC 1.99 12.77 CC 1.99
Shoulders MeanDose None 4.65 Gy 0.92 4.73 Gy 0.92
Final Score - - - 95.84 - 109.70
Table 10 and 11 shows the details scores and final score from the scorecard [32] to evaluate plan qualities for IMRT and VMAT, respectively. The structures (from the automatic pipeline) too small or with too much overlap with PTVs have been removed. For each table, we have included the RT structure names (i.e., StructId), Metric Type, the input for the Metric Type, output value, and the score of the output. We compare the output and its value of the plans from the initial (without the iterative refining) and full AIRTP, and we find that AIRTP generally improve the quality scores.
Table 11: An example of 4-arc VMAT for head-and-neck cancer.
StructId MetricType Input Output (Initial) Score (Initial) Output (AIRTP) Score (AIRTP)
PTVHighOPT VolumeAtDose 70.0 Gy 97.00 % 19.00 97.00 % 19.00
PTVHighOPT DoseAtVolume 99.5 % 69.00 Gy 1.32 69.02 Gy 1.33
PTVHighOPT DoseAtVolume 0.03 CC 74.39 Gy 8.86 74.19 Gy 9.01
PTVLow VolumeAtDose 56.0 Gy 97.21 % 14.53 97.79 % 14.63
PTVLow DoseAtVolume 99.5 % 54.51 Gy 1.17 55.11 Gy 1.30
PTVLow-PTVHigh VolumeAtDose 58.8 Gy 45.17 % 6.36 41.23 % 6.66
SpinalCord_05 DoseAtVolume 0.03 CC 41.38 Gy 6.09 38.60 Gy 6.16
SpinalCord_05 VolumeAtDose 40.0 Gy 0.49 % 1.95 0.00 % 2.00
SpinalCord_05 VolumeAtDose 30.0 Gy 33.79 % 1.12 10.67 % 1.82
BrainStem_03 DoseAtVolume 0.03 CC 45.32 Gy 3.12 44.47 Gy 3.14
Brain DoseAtVolume 0.03 CC 52.50 Gy 1.66 53.84 Gy 1.65
Brain VolumeAtDose 50.0 Gy 0.08 CC 3.00 0.11 CC 3.00
Pituitary MeanDose None 3.33 Gy 0.97 3.27 Gy 0.97
OpticNerve_L DoseAtVolume 0.03 CC 2.57 Gy 2.95 2.52 Gy 2.96
OpticNerve_R DoseAtVolume 0.03 CC 2.59 Gy 2.95 3.52 Gy 2.96
Cochlea_R VolumeAtDose 40.0 Gy 0.00 % 3.00 0.00 % 3.00
Cochlea_L VolumeAtDose 40.0 Gy 0.00 % 3.00 0.00 % 3.00
Lens_R DoseAtVolume 0.03 CC 1.54 Gy 2.35 1.41 Gy 2.36
Lips MeanDose None 23.99 Gy 3.01 17.38 Gy 5.79
Mandible-PTV VolumeAtDose 70.0 Gy 0.00 % 5.00 0.00 % 5.00
Mandible-PTV VolumeAtDose 60.0 Gy 0.00 % 2.00 0.00 % 2.00
Mandible-PTV VolumeAtDose 50.0 Gy 13.75 % 1.82 7.06 % 1.91
Esophagus MeanDose None 13.90 Gy 3.78 12.95 Gy 3.80
Esophagus DoseAtVolume 0.03 CC 48.52 Gy 3.19 53.59 Gy 3.06
OCavity-PTV MeanDose None 39.55 Gy 1.68 30.11 Gy 4.47
OCavity-PTV DoseAtVolume 0.03 CC 58.05 Gy 1.87 57.31 Gy 1.92
RingPTVHigh DoseAtVolume 0.03 CC 70.27 Gy 4.90 70.12 Gy 4.96
RingPTVLow DoseAtVolume 0.03 CC 58.23 Gy 1.52 57.95 Gy 2.29
Posterior_Neck VolumeAtDose 35.0 Gy 45.95 % 0.65 37.85 % 1.94
Trachea MeanDose None 10.84 Gy 2.17 8.31 Gy 2.25
Lungs VolumeAtDose 20.0 Gy 10.97 CC 1.99 9.92 CC 1.99
Shoulders MeanDose None 4.20 Gy 0.93 4.56 Gy 0.92
Final Score - - - 117.92 - 127.23
Appendix F Visualization of Creating Beam/Angle Plates

### v5

@@ Appendix B Challenges of Clinical Plans for AI Training
OAR 3694 Brain DNU, Brain Push, Brachial Plexus, Brain sub, Brain2, Brainstem expd, Bstem_PRV, Cord + 0.5cm, Cord<< 45, Cord_PRV, Cord+3, Cord+5mm
@@
As shown in Table 5, the naming of RT structures is highly heterogeneous. Naming conventions often vary across institutions, annotators, and even individual patients. Curating such datasets for AI training typically requires extensive manual review, despite the possibility of implementing some rule-based approaches. The presence of thousands of different — and sometimes subjective — structure names makes curation inefficient and severely limits the scalability of training data.
Fig. 8 further illustrates that the number of clinically contoured OARs varies significantly across patients, adding another layer of complexity for AI model development. In addition, the optimization objectives used to create clinical treatment plans are usually not recorded. These objectives are often complex, and their association with RT structures is not transparent, as some RT structures may not be included in the objectives.
While limited-scale datasets have enabled AI success in relatively simple scenarios — as referenced in Appendix A — they may require considerable effort in data curation. Building generalizable and robust AI models for real-world clinical settings remains a significant challenge. We believe that our work on preparing diverse clinical contexts and generating high-quality treatment plans at scale will be crucial for advancing the next generation of AI in radiotherapy planning.
Appendix C Standardized Structures and Score Card Items
Table 6: HaN RT structures for RapidPlan [magliari2024hn] and scorecard [varian2024han].
@@ Appendix C Standardized Structures and Score Card Items
Table 6 presents original RapidPlan [magliari2024hn] contouring items in head-and-neck cancer. Some organ contours may be missing due to several reasons: (1) the regions extend beyond the CT field of view, (2) there is significant overlap with clinical PTVs, or (3) segmentation failures occur, such as zero or extremely small segmented masks, (4) coustomized scorecard may not include all the structures. Additionally, as noted in the limitations, variations in contouring guidelines and individual planner preferences can introduce subjectivity. Simply combining clinical PTVs with auto-generated OAR masks and helper structures may not directly translate to clinical practice, as manual adjustments by planners are often necessary. While we cannot guarantee that every large-scale planning solution meets clinical standards, this approach still provides substantial value for AI model development, as noted in the main text.
@@
Since a publicly available RapidPlan model that adequately covers the various lung tumor locations is not found, we instead derive the optimization objectives directly from the scorecard for initial lung cancer planning, as shown in Table 8. The corresponding lung scorecard items are listed in Table 8.
Appendix D Data Curation Pipeline with AIRTP Integration
Figure 9: Overview of the data curation pipeline, from raw data processing to AI model training, highlighting where AIRTP integrates into the workflow.
The complete data curation pipeline is illustrated in Fig. 9. Raw data from public archives such as TCIA or from clinical collaborators are often not well organized. We use a DICOM parser tool to structure each study into separate folders containing CT, RTSTRUCT, RTPLAN (if available), and RTDOSE (if available) data. In Step 1 of the AIRTP pipeline, we segment all OARs, filter clinical PTVs through name and dose-level matching, and create helper structures, following the list provided in Appendix C. Along with the CT series, the newly generated RTSTRUCT data are batch-uploaded to the Eclipse server. Using Step 2 and 3 of AIRTP, we then generate high-quality RT plans at scale. Finally, DICOM files for CT, RTSTRUCT, RTPLAN, and RTDOSE are exported locally.
Exported DICOMs may be directly used for AI model training. Alternatively, researchers may process data and save them in more efficient formats (e.g., Numpy) to facilitate data loading. As an example, we provide a Python toolkit alongside the hosted GDP-HMM challenge to support data processing.
Appendix E Code Blocks for Algorithm 1
@@
Appendix F More about GDP-HMM Challenge
F.1 Data Distribution
@@
F.2 Baseline Method
As shown in Fig. 13, the educational baseline applies the MedNeXt as the backbone; due to its superior performance in multiple large-scale segmentation benchmarks. Fundamentally, the dose prediction is an conditional image generation problem. The CT images, PTV & OAR masks, planning mode, treatment site, and beam geometries etc., are conditions for 3D dose prediction. We follow the strategy of from [gao2023flexible], concatenating all the conditions as multi-channel 3D input. The loss function used in the baseline is pretty simple: the mean absolute error (MAE) of the predicted dose and the reference dose.
@@ F.2 Baseline Method
F.3 Result Samples from the Baseline
@@
We present four samples from our validation cohort, illustrating distinct scenarios that a universal AI model should be capable of handling. Fig. 14 and 15 depict treatment plans for the same head-and-neck cancer patient using different modalities: IMRT and VMAT. Similarly, Fig. 16 and 17 show treatment plans for the same lung cancer patient, also with IMRT and VMAT. Each case includes a comparison of the predicted and reference 3D dose distributions, along with corresponding dose-volume histograms (DVHs).
@@
Appendix G Visualization of Creating Beam/Angle Plates